# Error Correction in Dynamical Codes

### v1

@@ 5 Initialization of a Floquet Code
This maps $s_{1}\to s_{1},s_{2}\to s_{3}$ when $n=3$. To do this for the $4^{\mathrm{th}}$ cycle, we replace the variables in $M_{3}$ with the elements from the stabilizer group right before where the new sequence will be inserted using this mapping: $s_{1}\to d_{s_{1}}s_{2}s_{3}$, $s_{2}\to d_{s_{1}}s_{3}$, and $s_{3}\to s_{4}$. Using the updated variables, $d_{s_{1}}$ will be mapped to $d_{d_{s_{1}}s_{2}s_{3}}$ which anti-commutes with $d_{s_{1}}s_{2}s_{3}$ but commutes with all the other generators in the group including $d_{s_{1}}s_{2}$. Thus, the new sequence consists of measurements that commute with $d_{s_{1}}s_{2}$, so it will not be changed by the update from this set of new measurements, when it is inserted in the location indicated in Figure 7. This trick can be applied recursively for more stabilizers. To extend to $n$ stabilizers from $n-1$ stabilizers. Right before measuring the last measurement $s_{2}$, we replace $s_{1}$ with $d_{s_{1}}s_{2}s_{3},s_{2}$ with the stabilizer generator that is evolved from $s_{n-1}$, and replace $s_{3}$ with $s_{n}$. One can verify that inserting the new measurement sequence does not affect the previous cycles. Since $\langle S_{j}^{(i)}\rangle\subseteq\langle S_{j+1}^{(i)}\rangle$, the stabilizer generators before making the measurements in the new sequence are a subset of the current cycle, so the new sequence will act as identity to those generators. One can also verify that in the future cycles, $s_{1}$ will still be mapped to $s_{1}$ and $s_{i}\to s_{i+1}$ because the new sequence maps $s_{n-1}\to s_{n}$ while preserving the other mappings. ∎
@@ 6 Geometrically local dynamical code
The honeycomb code and other generalizations of Floquet codes are geometrically local codes.
@@
[24] If the generators of the gauge group $\mathcal{G}$ are geometrically local, then the distance satisfies $d\leq 3rL^{D-1}$, where $r$ is the width of the local gauges. In particular, local stabilizer codes have distance given by $O(L^{D-1})$.
@@
[30] Local stabilizer codes satisfy the following constrain:
@@ 6.2 Geometrically local stabilizer codes with long range stabilizers
First we split the D dimensional lattice into unit cells with volume $v_{c}=\alpha d$, where $d$ is the distance and $\alpha$ is a constant. For a fixed constant $r=O(1)$, we can pick a ball of radius $r$ from each cell, such that each ball is spatially disjoint from the other balls and from the boundaries of the lattice by a constant physical distance that is at least $r_{\mathrm{local}}$ (See Figure 8). Let $R_{0}^{\prime}$ denote the union of the balls. $R_{0}^{\prime}$ may have overlap with $H^{\prime}:=\mathcal{B}(Q,(2^{D}+1)s_{U}+c^{\prime})$, where $c^{\prime}>2r_{\mathrm{local}}$. We deform $H^{\prime}$ locally so that the balls that overlap with $H^{\prime}$ are now contained in $H^{\prime}$. Let $c_{0}$ denote the maximum physical distance of the deformation from the boundary of $H^{\prime}$. Then, we set $c=c_{0}+c^{\prime}$ and define $H_{1}$ as $H_{1}:=\mathcal{B}(Q,(2^{D}+1)s_{U}+c)$, with $H^{\prime}\subseteq H_{1}$. Let $R_{0}:=R_{0}^{\prime}\backslash H^{\prime}$. This is the set of balls that are far away from $H^{\prime}$. We want to show that $R_{0}$ forms a bare cleanable region. Since each ball is $o(d)$, they are each bare cleanable. Further, we have picked the balls to be supported away from the boundaries of the unit cells, such that the balls are at least $2r_{\mathrm{local}}$ apart, so any local stabilizer generator overlaps with at most 1 ball and since $R_{0}$ does not overlap with $H^{\prime}$, any long range stabilizers that are only supported on $Q$ are spatially disjoint from $R_{0}$. Thus, the Union Lemma can be applied to $R_{0}$, and we conclude that $R_{0}$ forms a bare cleanable region. Using $R_{0}^{\prime}$ as the set of mutually disjoint balls on the code, one can draw lines connecting these balls, then fatten the lines to form a covering for all the physical qubits, following the same construction in PY’s result. This gives a covering of the full lattice with $R_{m}^{\prime}$ for $m\in[0,D]$. $R_{m}^{\prime}$ consists of $m$ dimensional connected components. $R_{D}^{\prime}$ consists of $D$ dimensional skewed cells. See Figure 8 for illustration.
@@
As the new regions can overlap with $H^{\prime}$, we define the regions as follow: $R_{m}:=R_{m}^{\prime}\backslash H^{\prime}$, for $1\leq m\leq D$. We show that each region $R_{m}$ is bare cleanable. The volume of each connected component $r_{m}^{\prime}$ of $R_{m}^{\prime}$ is at most $O(\alpha d^{\frac{m}{D}})=O(d)$. Since any region with volume less than the distance $d=\Omega(\log(n))$ is bare cleanable, $r_{m}^{\prime}\backslash H^{\prime}=O(d)$ is also a bare cleanable object. Since each $r_{m}^{\prime}$ is spatially disjoint, and each $r_{m}^{\prime}\backslash H^{\prime}$ is also spatially disjoint. Supported away from the qubits with long range stabilizers, $r_{m}^{\prime}\backslash H^{\prime}$ can also be cleaned using local stabilizers that have no overlap with other $m$-dimensional objects. Thus, we can apply the Union Lemma to conclude that $R_{m}$ is bare cleanable. Given a constant depth circuit, we have to check that each region $R_{m}^{+}:=\mathcal{B}(R_{m},2^{m-1}s_{U})$ is still bare cleanable for $m>0$, which is a condition required for the next part of this proof. For each connected component $r_{m}^{\prime}$ in $R_{m}^{\prime}$, any Pauli operator supported on $r_{m}^{\prime}\backslash H^{\prime}$ is spread locally since it is at least $(2^{D}+1)s_{U}$ away from $Q$, the region that does support long range connectivity. Further, $\mathcal{B}(r_{m}^{\prime}\backslash H^{\prime},2^{m-1}s_{U})$ has no overlap with $Q$, so cleaning $\mathcal{B}(r_{m}^{\prime}\backslash H^{\prime},2^{m-1}s_{U})$ does not spread Pauli operators supported on one of the $r_{m}^{\prime}\backslash H^{\prime}$ to other objects in $R_{m}^{+}$ by any long range stabilizers. Since $r_{m}^{\prime}\backslash H^{\prime}$ increases by a constant radius of $2^{m-1}s_{U}=O(1)$, $\mathcal{B}(r_{m}^{\prime}\backslash H^{\prime},2^{m-1}s_{U})$ remains bare cleanable. By choosing $r$ to be greater than $2^{m}s_{U}$, each object of the form $\mathcal{B}(r_{m}^{\prime}\backslash H^{\prime},2^{m-1}s_{U})$ is disconnected from the other objects in $R_{m}^{+}$. Therefore, the union of $R_{m}^{+}$ is still bare cleanable.
@@
Next we prove that $U$ must be in $\mathcal{C}^{(D)}$, with respect to the logical qubit $q_{L}$. Suppose the logical representatives $\{x_{L},z_{L}\}$ are fully supported on $\overline{H}:=\overline{\mathcal{B}(H_{1},2^{D}s_{U})}$. The proof holds if we replace $H_{1}$ with $H^{\prime}$ but for simplicity and better clarity we will use $H_{1}$ instead. Without assuming that $H$ is bare cleanable, we pick two arbitrary logical representatives $L_{1}$ and $L_{2}$ for the logical qubit $q_{L}$ such that they are fully supported on $\overline{\mathcal{B}(H_{1},2^{D}s_{U})}$. We show that $U\in\mathcal{C}^{(2)}$ with respect to the logical qubit for $D=2$ before generalizing the proof to arbitrary dimension. For $D=2$, $H_{1}=\mathcal{B}(Q,5s_{U}+c),$ $R_{1}^{+}=\mathcal{B}(R_{1},s_{U})$, $R_{2}^{+}=\mathcal{B}(R_{2},2s_{U})$ $L_{1}$ can be cleaned from $R_{1}^{+}$, since $R_{1}^{+}$ is dressed cleanable. After cleaning, $L_{1}$ is supported on:
@@
Let $L_{1}^{\prime}:=UL_{1}U^{\dagger}L_{1}^{\dagger}$ and $L_{2}^{\prime}=L_{2}L_{1}^{\prime}L_{2}^{\dagger}L_{1}^{\prime\dagger}$. We show that $L_{2}^{\prime}$ is an identity operator up to ±$\pm 1$ phase. This will imply that $L_{1}^{\prime}$ is a Pauli operator so $U$ is at most a Clifford gate on $q_{L}$.
@@
We find that $L_{1}^{\prime}$ is fully supported on $R_{0},R_{2}$ and $\mathcal{B}(\partial\mathcal{B}(Q,4s_{U}+c)\cap\partial_{+}R_{1}^{+},s_{U})$. $L_{2}^{\prime}$ is fully supported on the region that supports $L_{1}^{\prime}$, by treating $L_{2}$ as a unitary operator acting on $L_{1}^{\prime}$. Treating $L_{1}^{\prime}$ as a unitary operator acting on $L_{2}$, we also obtain that $L_{2}^{\prime}$ is fully supported on $\mathcal{B}(D_{2},2s_{U})$, which is given by:
@@
First two terms are in $R_{0}$ and $R_{1}$ respectively. The last term can be rewritten as $\partial\mathcal{B}(Q,3s_{U}+c)\cap\partial_{+}R_{2}^{+}$, with a spread of constant $2s_{U}$ which is local since the support is at least $s_{U}$ away from $Q$. Thus, $L_{2}^{\prime}$ must be supported on $R_{0}$ and the intersection between the two different boundaries each with some constant spread. However this is upper bounded by the support on the union of $R_{0}$ and the overlap of $R_{1}^{+}$ and $R_{2}^{+}$ with a constant local spread of $3s_{U}$. These boundary objects have volume at most $O(d^{\frac{D-1}{D}})$ by construction, since they have a constant width in one of the dimensions. Thus, each object is bare cleanable, and by construction, they are spatially disjoint objects. Further, they are supported only on the boundary of $H_{1}$, so they can be cleaned by local stabilizers. Thus, by the Union Lemma, they form a bare cleanable set. As the boundary objects have no overlap with $R_{0}$ since they must be fully supported on the boundary of $\mathcal{B}(Q,4s_{U}+c)$ and $\mathcal{B}(Q,3s_{U}+c)$ with a constant spread of $2s_{U}$ at most, which is disjoint from $R_{0}$ by construction, the union of $R_{0}$ and the boundaries is also bare cleanable. Hence, $L_{2}^{\prime}$ is an identity operator. This implies that $L_{1}^{\prime}$ is a Pauli operator so $U$ is at most a Clifford gate on $q_{L}$.

### v2

@@ 6 Initialization of a Floquet Code
This maps $s_{1}\to s_{1},s_{2}\to s_{3}$ when $n=3$. To do this for the $4^{\mathrm{th}}$ cycle, we replace the variables in $M_{3}$ with the elements from the stabilizer group right before where the new sequence will be inserted using this mapping: $s_{1}\to d_{s_{1}}s_{2}s_{3}$, $s_{2}\to d_{s_{1}}s_{3}$, and $s_{3}\to s_{4}$. Using the updated variables, $d_{s_{1}}$ will be mapped to $d_{d_{s_{1}}s_{2}s_{3}}$ which anti-commutes with $d_{s_{1}}s_{2}s_{3}$ but commutes with all the other generators in the group including $d_{s_{1}}s_{2}$. Thus, the new sequence consists of measurements that commute with $d_{s_{1}}s_{2}$, so it will not be changed by the update from this set of new measurements, when it is inserted in the location indicated in Figure 7.
This trick can be applied recursively for more stabilizers. To extend to $n$ stabilizers from n1n-1 stabilizers. Right before measuring the last measurement $s_{2}$, we replace $s_{1}$ with $d_{s_{1}}s_{2}s_{3},s_{2}$ with the stabilizer generator that is evolved from $s_{n-1}$, and replace $s_{3}$ with $s_{n}$.
One can verify that inserting the new measurement sequence does not affect the previous cycles. Since $\langle S_{j}^{(i)}\rangle\subseteq\langle S_{j+1}^{(i)}\rangle$, the stabilizer generators before making the measurements in the new sequence are a subset of the current cycle, so the new sequence will act as identity to those generators. One can also verify that in the future cycles, $s_{1}$ will still be mapped to $s_{1}$ and $s_{i}\to s_{i+1}$ because the new sequence maps $s_{n-1}\to s_{n}$ while preserving the other mappings. ∎
Example 50.
We provide an alternative construction below that looks deceivingly simple, using constant cycle size and also purely local measurements (in 1D), yet takes $O(n)$ time to initialize/learn the syndrome:
Round 1) Measure $X_{1}$
Round 2) Measure $X_{2}X_{3}$, $X_{6}X_{7}$, and so on ($X_{4i+2}X_{4i+3}$)
Round 3) Measure $Z_{1}Z_{2}$, $Z_{5}Z_{6}$, and so on ($Z_{4i+1}Z_{4i+2}$)
Round 4) Measure $X_{4}X_{5}$, $X_{8}X_{9}$, and so on ($X_{4i}X_{4i+1}$)
Round 5) Measure $Z_{3}Z_{4}$, $Z_{7}Z_{8}$, and so on ($Z_{4i+3}Z_{4i+4}$)
Round $k>5$) Repeat Round $((k-2)\mod 4)+2$.
We explicitly show the ISGs for $n=10$, in which case we have round 1, followed by cycles of 4 rounds of measurements given by: $\mathrm{Round\;1}:\mathrm{Measure\;}X_{2}X_{3},X_{6}X_{7}$. $\mathrm{Round\;2}:\mathrm{Measure\;}Z_{1}Z_{2},Z_{5}Z_{6},Z_{9}Z_{10}$. $\mathrm{Round\;3}:\mathrm{Measure\;}X_{4}X_{5},X_{8}X_{9}$. $\mathrm{Round\;4}:\mathrm{Measure\;}Z_{3}Z_{4},Z_{7}Z_{8}$
$\mathrm{Round\;1)}$ $X_{1}$
$\mathrm{Round\;2)}$ $X_{1},X_{2}X_{3},X_{6}X_{7},$
$\mathrm{Round\;3)}$ $Z_{1}Z_{2},Z_{5}Z_{6},Z_{9}Z_{10},X_{1}X_{2}X_{3}$
$\mathrm{Round\;4)}$ $X_{1}X_{2}X_{3},Z_{1}Z_{2},X_{4}X_{5},X_{8}X_{9}$
$\mathrm{Round\;5)}$ $Z_{1}Z_{2},Z_{3}Z_{4},Z_{7}Z_{8},X_{1}X_{2}X_{3}X_{4}X_{5}$
$\mathrm{Round\;6)}$ $X_{1}X_{2}X_{3}X_{4}X_{5},X_{2}X_{3},X_{6}X_{7},Z_{1}Z_{2}Z_{3}Z_{4}$
$\mathrm{Round\;7)}$ $X_{1}X_{2}X_{3}X_{4}X_{5}X_{6}X_{7},Z_{1}Z_{2}Z_{3}Z_{4},Z_{1}Z_{2},Z_{5}Z_{6},Z_{9}Z_{10}$
$\mathrm{Round\;8)}$ $X_{1}X_{2}X_{3}X_{4}X_{5}X_{6}X_{7},Z_{1}Z_{2}Z_{3}Z_{4}Z_{5}Z_{6},X_{4}X_{5},X_{8}X_{9},Z_{1}Z_{2}$
$\mathrm{Round\;9)}$ $X_{1}X_{2}X_{3}X_{4}X_{5}X_{6}X_{7}X_{8}X_{9},Z_{1}Z_{2}Z_{3}Z_{4}Z_{5}Z_{6},Z_{3}Z_{4},Z_{7}Z_{8},Z_{1}Z_{2}$
$\mathrm{Round\;10)}$ $X_{1}X_{2}X_{3}X_{4}X_{5}X_{6}X_{7}X_{8}X_{9},Z_{1}Z_{2}Z_{3}Z_{4}Z_{5}Z_{6}Z_{7}Z_{8},X_{6}X_{7},Z_{3}Z_{4},Z_{1}Z_{2}$
$\mathrm{Round\;11)}$ $X_{1}X_{2}X_{3}X_{4}X_{5}X_{6}X_{7}X_{8}X_{9},Z_{1}Z_{2}Z_{3}Z_{4}Z_{5}Z_{6}Z_{7}Z_{8},X_{6}X_{7},X_{2}X_{3},Z_{1}Z_{2}Z_{3}Z_{4}$
$\mathrm{Round\;12)}$ $X_{1}X_{2}X_{3}X_{4}X_{5}X_{6}X_{7}X_{8}X_{9},Z_{1}Z_{2}Z_{3}Z_{4}Z_{5}Z_{6}Z_{7}Z_{8},Z_{1}Z_{2}Z_{3}Z_{4},Z_{1}Z_{2},Z_{5}Z_{6},Z_{9}Z_{10}$
$\mathrm{Round\;13)}$ $X_{1}X_{2}X_{3}X_{4}X_{5}X_{6}X_{7}X_{8}X_{9},Z_{1}Z_{2}Z_{3}Z_{4}Z_{5}Z_{6}Z_{7}Z_{8}Z_{9}Z_{10},Z_{1}Z_{2}Z_{3}Z_{4}Z_{5}Z_{6},X_{4}X_{5},X_{8}X_{9},Z_{1}Z_{2}$
Notice that for the chain of $X_{1}X_{2}X_{3}X_{4}\cdots X_{n}$ to form, it takes $n$ rounds. Thus, the initialization requires at least (n1)/4=O(n)(n-1)/4=O(n) cycles.
@@ 7 Geometrically local dynamical code
The honeycomb code and other generalizations of Floquet codes are geometrically local codes. In this section, we consider some properties of dynamical codes (not necessarily Floquet codes with repeating cycles) for which the measurements are all geometrically local.
@@
[28]
If the generators of the gauge group $\mathcal{G}$ are geometrically local, then the distance satisfies $d\leq 3rL^{D-1}$, where $r$ is the width of the local gauges. In particular, local stabilizer codes have distance given by $O(L^{D-1})$.
@@
[34]
Local stabilizer codes satisfy the following constrain:
@@ 7.2 Geometrically local stabilizer codes with long range stabilizers
First we split the D dimensional lattice into unit cells with volume $v_{c}=\alpha d$, where $d$ is the distance and $\alpha$ is a constant. For a fixed constant $r=O(1)$, we can pick a ball of radius $r$ from each cell, such that each ball is spatially disjoint from the other balls and from the boundaries of the lattice by a constant physical distance that is at least $r_{\mathrm{local}}$ (See Figure 8).
Let $R_{0}^{\prime}$ denote the union of the balls. $R_{0}^{\prime}$ may have overlap with $H^{\prime}:=\mathcal{B}(Q,(2^{D}+1)s_{U}+c^{\prime})$, where $c^{\prime}>2r_{\mathrm{local}}$. We deform $H^{\prime}$ locally so that the balls that overlap with $H^{\prime}$ are now contained in $H^{\prime}$. Let $c_{0}$ denote the maximum physical distance of the deformation from the boundary of $H^{\prime}$. Then, we set $c=c_{0}+c^{\prime}$ and define $H_{1}$ as $H_{1}:=\mathcal{B}(Q,(2^{D}+1)s_{U}+c)$, with $H^{\prime}\subseteq H_{1}$.
[…]
Using $R_{0}^{\prime}$ as the set of mutually disjoint balls on the code, one can draw lines connecting these balls, then fatten the lines to form a covering for all the physical qubits, following the same construction in PY’s result. This gives a covering of the full lattice with $R_{m}^{\prime}$ for $m\in[0,D]$. $R_{m}^{\prime}$ consists of $m$ dimensional connected components. $R_{D}^{\prime}$ consists of $D$ dimensional skewed cells. See Figure 8 for illustration.
@@
[…]
Given a constant depth circuit, we have to check that each region $R_{m}^{+}:=\mathcal{B}(R_{m},2^{m-1}s_{U})$ is still bare cleanable for $m>0$, which is a condition required for the next part of this proof. For each connected component $r_{m}^{\prime}$ in $R_{m}^{\prime}$, any Pauli operator supported on $r_{m}^{\prime}\backslash H^{\prime}$ is spread locally since it is at least $(2^{D}+1)s_{U}$ away from $Q$, the region that does support long range connectivity. Further, $\mathcal{B}(r_{m}^{\prime}\backslash H^{\prime},2^{m-1}s_{U})$ has no overlap with $Q$, so cleaning $\mathcal{B}(r_{m}^{\prime}\backslash H^{\prime},2^{m-1}s_{U})$ does not spread Pauli operators supported on one of the $r_{m}^{\prime}\backslash H^{\prime}$ to other objects in $R_{m}^{+}$ by any long range stabilizers. Since $r_{m}^{\prime}\backslash H^{\prime}$ increases by a constant radius of $2^{m-1}s_{U}=O(1)$, $\mathcal{B}(r_{m}^{\prime}\backslash H^{\prime},2^{m-1}s_{U})$ remains bare cleanable. By choosing $r$ to be greater than $2^{m}s_{U}$, each object of the form $\mathcal{B}(r_{m}^{\prime}\backslash H^{\prime},2^{m-1}s_{U})$ is disconnected from the other objects in $R_{m}^{+}$. Therefore, the union of $R_{m}^{+}$ is still bare cleanable.
@@
Next we prove that $U$ must be in $\mathcal{C}^{(D)}$, with respect to the logical qubit $q_{L}$. Suppose the logical representatives $\{x_{L},z_{L}\}$ are fully supported on $\overline{H}:=\overline{\mathcal{B}(H_{1},2^{D}s_{U})}$. The proof holds if we replace $H_{1}$ with $H^{\prime}$ but for simplicity and better clarity we will use $H_{1}$ instead.
Without assuming that $H$ is bare cleanable, we pick two arbitrary logical representatives $L_{1}$ and $L_{2}$ for the logical qubit $q_{L}$ such that they are fully supported on $\overline{\mathcal{B}(H_{1},2^{D}s_{U})}$. We show that $U\in\mathcal{C}^{(2)}$ with respect to the logical qubit for $D=2$ before generalizing the proof to arbitrary dimension. For $D=2$, $H_{1}=\mathcal{B}(Q,5s_{U}+c),$ $R_{1}^{+}=\mathcal{B}(R_{1},s_{U})$, $R_{2}^{+}=\mathcal{B}(R_{2},2s_{U})$
$L_{1}$ can be cleaned from $R_{1}^{+}$, since $R_{1}^{+}$ is dressed cleanable. After cleaning, $L_{1}$ is supported on:
@@
Let $L_{1}^{\prime}:=UL_{1}U^{\dagger}L_{1}^{\dagger}$ and $L_{2}^{\prime}=L_{2}L_{1}^{\prime}L_{2}^{\dagger}L_{1}^{\prime\dagger}$. We show that $L_{2}^{\prime}$ is an identity operator up to $\pm 1$ phase. This will imply that $L_{1}^{\prime}$ is a Pauli operator so $U$ is at most a Clifford gate on $q_{L}$.
@@
We find that $L_{1}^{\prime}$ is fully supported on $R_{0},R_{2}$ and $\mathcal{B}(\partial\mathcal{B}(Q,4s_{U}+c)\cap\partial_{+}R_{1}^{+},s_{U})$.
$L_{2}^{\prime}$ is fully supported on the region that supports $L_{1}^{\prime}$, by treating $L_{2}$ as a unitary operator acting on $L_{1}^{\prime}$. Treating $L_{1}^{\prime}$ as a unitary operator acting on $L_{2}$, we also obtain that $L_{2}^{\prime}$ is fully supported on $\mathcal{B}(D_{2},2s_{U})$, which is given by:
@@
First two terms are in $R_{0}$ and $R_{1}$ respectively. The last term can be rewritten as $\partial\mathcal{B}(Q,3s_{U}+c)\cap\partial_{+}R_{2}^{+}$, with a spread of constant $2s_{U}$ which is local since the support is at least $s_{U}$ away from $Q$.
Thus, $L_{2}^{\prime}$ must be supported on $R_{0}$ and the intersection between the two different boundaries each with some constant spread. However this is upper bounded by the support on the union of $R_{0}$ and the overlap of $R_{1}^{+}$ and $R_{2}^{+}$ with a constant local spread of $3s_{U}$. These boundary objects have volume at most $O(d^{\frac{D-1}{D}})$ by construction, since they have a constant width in one of the dimensions. Thus, each object is bare cleanable, and by construction, they are spatially disjoint objects. Further, they are supported only on the boundary of $H_{1}$, so they can be cleaned by local stabilizers. Thus, by the Union Lemma, they form a bare cleanable set. As the boundary objects have no overlap with $R_{0}$ since they must be fully supported on the boundary of $\mathcal{B}(Q,4s_{U}+c)$ and $\mathcal{B}(Q,3s_{U}+c)$ with a constant spread of $2s_{U}$ at most, which is disjoint from $R_{0}$ by construction, the union of $R_{0}$ and the boundaries is also bare cleanable. Hence, $L_{2}^{\prime}$ is an identity operator. This implies that $L_{1}^{\prime}$ is a Pauli operator so $U$ is at most a Clifford gate on $q_{L}$.